Littleton, Inc., has fixed costs of $75,000 per month, variable costs of $5 per unit, and a sales price per unit of $30. What is the break-even quantity per month
Answer #1

Contribution margin=Sales-Variable costs

=(30-5)=$25 per unit

Hence breakeven quantity=Fixed cost/Contribution margin

=(75000/25)

=3000 units per month.